MOA-II Obsevation Log Monthly Summary 2007/3
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 5042-5590 | 549 | 2660 | 150900 |
LMC | 1308-1401 | 94 | 511 | 136020 |
SMC | 342-350 | 9 | 17 | 4980 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 652 | 3188 | 291900 (81h05m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
99 | 98 | 97 | 269 | 430 | 83 | 92 | 85 | 425 | 104 | 83 | 80 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
80 | 74 | 74 | 77 | 72 | 71 | 70 | 67 | 63 | 60 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
20 | 20 | 20 | 16 | 16 | 22 | 94 | 90 | 88 | 27 | 22 | 23 |
Field13 | Field14 | Field15 | Field16 |
26 | 27 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
8 | 9 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2007/3/1: D2139(300s) / night time = 31560(sec)
2007/3/2: D2140(60s) / night time = 31800(sec)
2007/3/3: D2141 / night time = 31980(sec)
2007/3/4: Camera vacuum sensor failure. D2142(180) D2143(60) / night time = 32220(sec)
2007/3/5: D1244(240) D1245(180) D1246(60) The seeing was bad tonight / night time = 32460(sec)
2007/3/6: Overcast all night / night time = 32640(sec)
2007/3/7: / night time = 32940(sec)
2007/3/8: D2147(300s), D2148(60s) / night time = 33180(sec)
2007/3/9: D2149(300s) , D2150(60s) / night time = 33360(sec)
2007/3/10: D2151(60s), D2152(300s) / night time = 33600(sec)
2007/3/11: D2153 and D2154 300s darks have a strange patern which may be due to high camera temperature (-55) / night time = 33780(sec)
2007/3/12: No observing due to bad weather / night time = 34020(sec)
2007/3/13: Lots of condensation. D2155(60s) / night time = 34200(sec)
2007/3/14: D2156 (60s) , D2157 (300s) / night time = 34440(sec)
2007/3/15: No observing due to bad weather / night time = 34620(sec)
2007/3/16: No observing due to bad weather / night time = 34860(sec)
2007/3/17: No observing due to bad weather / night time = 31440(sec)
2007/3/18: No observing due to strong wind / night time = 35280(sec)
2007/3/19: D2164 / night time = 35460(sec)
2007/3/20: This online log was not recorded by Paul. D2165(300sec), D2166(60sec) are OK. / night time = 35640(sec)
2007/3/21: D2168(300), D2167(60) / night time = 35880(sec)
2007/3/22: Warm camera, Camera Error at start of night. Tel driver unit error at start(Zero search). D2169(60), D2170(300) / night time = 36060(sec)
2007/3/23: Cam warm at 7:30 (WAITING). D2171(60), D2172(300) / night time = 36240(sec)
2007/3/24: 2173(60), 2174(300) / night time = 36480(sec)
2007/3/25: 2175(60), 2176(300) / night time = 36660(sec)
2007/3/26: Thick cloud untill 4:10 then some observing to get alert event. (D2177 60s) / night time = 36840(sec)
2007/3/27: D2178(60), D2179(300) / night time = 37080(sec)
2007/3/28: This online log was not recorded By Paul. No observation / night time = 37260(sec)
2007/3/29: D2180(60), / night time = 37440(sec)
2007/3/30: D2181(60), / night time = 37680(sec)
2007/3/31: No Observing / night time = 37800(sec)
TOTAL NIGHT TIME: 1074900(sec)
OPERATING RATE: 39.0 %
GB: 23.9 %
LMC: 14.6 %
SMC: 0.5 %
ALERT: 0.0 %
GRB: 0.0 %
Return to index
Previous month
Next month